OpenStudy (anonymous):

4 atm is the partial pressure of argon in a mixture.there is 1 mole of H2 .3.5 mole of He ,1.5 mole of N2 & 3 mole argon.find total pressure...

OpenStudy (anonymous):

Use Dalton's law of partial pressures:\[P_t=(n_1+n_2+n_3+...+n_i)(RT/V)\] http://en.wikipedia.org/wiki/Dalton's_law

OpenStudy (anonymous):

In this case, Dalton's law is:\[P_{total}=P_{H}+P_{He}+P_{N}+P_{Ar}=(n_{H}+n_{He}+n_{N}+n_{Ar})(RT/V)\] We don't know what T or V is, but we can solve for (RT/V) by plugging in the values for Ar.\[P_{Ar}=(n_{Ar})(RT/V)\]\[(RT/V)=(P_{Ar})/(n_{Ar})=(4/3)(atm/mol)\] Then you can solve for total pressure by using the supplied values and the calculated value of (RT/V): \[P_{total}=(n_{H}+n_{He}+n_{N}+n_{Ar})(RT/V)=(1.0+3.5+1.5+3.0)(4/3)(atm)\]\[P_{total}=(9.0)(4/3)(atm)=12atm\]

OpenStudy (anonymous):

thank you... :)

OpenStudy (anonymous):

Very welcome.